# Foundational Automatic Evaluators: Scaling Multi-Task Generative Evaluator Training for Reasoning-Centric Domains

### v1

@@ D.2 How does direct judgment prompting affect performance?
A known issue in pairwise evaluation is inconsistency (Wang et al., 2023b), a form of positional bias where the evaluator judgment changes based on the order of responses in the input prompt. During training, we observed that our judges become more consistent as a function of data scale; Fig. 6 shows the progression of pairwise consistency on the five pairwise benchmarks in § 4.1 over the course of training for FARE-8B and an earlier training run which was initialized from Qwen2.5-7B-Instruct. Both models steadily become positionally robust over the course of training, with the weaker Qwen2.5-7B-Instruct showing substantial gains. This reveals that scaling evaluator training data can mitigate common judge biases, complementing mitigation strategies that use data augmentation (Saha et al., 2025), label balancing (Cao et al., 2024; Wang et al., 2024a), and RL-based reward or algorithmic methods (Whitehouse et al., 2025; Xu et al., 2025b).
@@
We additionally evaluate FARE on Single Rating tasks with BiGGen-Bench (Kim et al., 2024a) and FLASK (Ye et al., 2023), two chat-centric evaluation datasets with human annotated 1-5 ratings. We measure Pearson correlation with human annotations, and report results in Table 8. Single-rating is widely used evaluation task in reasoning settings, and thus constituting the smallest proportion of our training data, as shown in Fig. 2. Nonetheless, FARE are competitive with chat-focused judge models, with FARE-8B outperforming foundational judge models like SFR-Judge-8B and 12B and FARE-20B approaching the performance of SFR-Judge-70B. We use reported values from baseline papers, including LMUnit (Saad-Falcon et al., 2024), Atla Selene (Alexandru et al., 2025), and SFR-Judge (Wang et al., 2024a).
Table 9: Comparison of FARE against their initial models and other popular general-purpose models. indicates some results reported in Whitehouse et al. (2025) or Zheng et al. (2024).
@@ E.2 Sample evaluation rubric
Here, we provide a sample rubric that was hand-written for SWE-Rank (Reddy et al., 2025). SWE-Rank data consists of contrastive pairs for training retrieval models. We re-purposed this data into a binary verification task, asking the evaluator if the retrieved code snippet was relevant for editing given a user request. “Positive” samples were assigned “Correct” labels, and “Negative” samples were assigned “Incorrect” labels.

### v2

@@ D.2 How does direct judgment prompting affect performance?
A known issue in pairwise evaluation is inconsistency (wang2023large), a form of positional bias where the evaluator judgment changes based on the order of responses in the input prompt. During training, we observed that our judges become more consistent as a function of data scale; Fig. 6 shows the progression of pairwise consistency on the five pairwise benchmarks in § 4.1 over the course of training for FARE-8B and an earlier training run which was initialized from Qwen2.5-7B-Instruct. Both models steadily become positionally robust over the course of training, with the weaker Qwen2.5-7B-Instruct showing substantial gains. This reveals that scaling evaluator training data can mitigate common judge biases, complementing mitigation strategies that use data augmentation (saha2025learning), label balancing (cao2024compassjudger; wang2024direct), and RL-based reward or algorithmic methods (whitehouse2025j1; xu2025j4r).
@@
We additionally evaluate FARE on Single Rating tasks with BiGGen-Bench (kim2024biggen) and FLASK (ye2023flask), two chat-centric evaluation datasets with human annotated 1-5 ratings. We measure Pearson correlation with human annotations, and report results in Table 8. Single-rating is widely used evaluation task in reasoning settings, and thus constituting the smallest proportion of our training data, as shown in Fig. 2. Nonetheless, FARE are competitive with chat-focused judge models, with FARE-8B outperforming foundational judge models like SFR-Judge-8B and 12B and FARE-20B approaching the performance of SFR-Judge-70B. We use reported values from baseline papers, including LMUnit (saad2024lmunit), Atla Selene (alexandru2025atla), and SFR-Judge (wang2024direct).
Table 9: Comparison of FARE against their initial models and other popular general-purpose models. indicates some results reported in whitehouse2025j1 or zheng2024processbench.
@@ E.2 Sample evaluation rubric
Here, we provide a sample rubric that was hand-written for SWE-Rank (reddy2025swerank). SWE-Rank data consists of contrastive pairs for training retrieval models. We re-purposed this data into a binary verification task, asking the evaluator if the retrieved code snippet was relevant for editing given a user request. “Positive” samples were assigned “Correct” labels, and “Negative” samples were assigned “Incorrect” labels.